How to find the weight of element in a given weight of compund?


For eg. Find the weight of Carbon in 99 g of C02

Grade:10

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

44 gm CO2 has weight 12 gm carbon.

Then weight of 99gm CO2=(12/44)*99

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

99gm CO2 = 99/44 mol of CO2                    (molar mass of CO2 = 44)

1 mol CO2  gives 1mol carbon + 2 mol oxygen

99/44 mol CO2 gives 99/44 mol carbon + 2(99/44) mol oxygen

 99/44 mol carbon = 12(99/44) gram

                          =27 gram C

aatish agrawal
13 years ago

i mole of co2 =44g ,it contains 12 g of carbon

so 99g of co2 will contain (99*12)/44 =27 g ...